OpenStudy (anonymous):

"The perimeter of a rectangle is 18 inches. The length is three more than the width. What is the length?" So far I've got: P=18 L+3=W And so I think [(L+3)+w]2=P But I could be wrong...Any takers?

OpenStudy (anonymous):

P = 2(L+W)

OpenStudy (anonymous):

If you want just one variable (and I think you do) w = L+w means that L = w-3 using daniel's equation P=2((w-3)+w) ya? or 18=2w-6 and then it is easy

OpenStudy (anonymous):

2(l+w) = 18 2(w+3+w)=18 since l is w+3 substitute .. 2w+3=9 w=3

OpenStudy (anonymous):

aaak that should be w = L +3 but you knew that

OpenStudy (anonymous):

w=3 so l is 6

OpenStudy (toxicsugar22):

hi cna u help me after her